OpenStudy (anonymous):

Terrell and Mindi are organizing a neighborhood skateboarding competition and need to build a mini ramp. Terrell can build the ramp in 12 hours. Mindi can build the ramp in 6 hours. Which shows the equation and solution for the time it will it take Terrell and Mindi to build the ramp if they work together? Round your answer to the nearest hour, if necessary.

hero (hero):

I have an equation and solution but it may not match yours

OpenStudy (anonymous):

they are both losers because i can build a ramp in ten minutes \

OpenStudy (anonymous):

if it will help then give it a shot

OpenStudy (anonymous):

ok well that was no help bigc84

hero (hero):

\[\frac{12 \times 6}{12 + 6} = x\]

OpenStudy (anonymous):

well ill give you the options hero

hero (hero):

\(x =\) Time it takes both of them while working together

hero (hero):

I'm not going to give you the answer. You have to take my hint and run with it.

OpenStudy (anonymous):

i know. hahaXD @ineedhelp72 Thats what i was thinkin @Hero

OpenStudy (anonymous):

i dont want the answer i want to know how to get each part of the answer

hero (hero):

Start by solving for x in the equation I gave you.

OpenStudy (anonymous):

A.2=(1/12+1/6)t; 3hours B.2=(12/1+6/1)t; 4hours C.1=(1/12+1/6)t; 4hours D.1=(12/1+6/1)t; 3hours

OpenStudy (anonymous):

i know trhe anser is going to be (1/12+1/6) but i dont know how to get the hours and what t; 3 or 4 means?

hero (hero):

Neveermind

hero (hero):

But anyway, you should try solving the equations

OpenStudy (anonymous):

which ones?